If 8,000 pesos is the present value of 11,000 pesos due at the end of seven months, what is the simple discount?


1
Expert's answer
2021-09-29T11:29:54-0400
"\frac{11000-8000}{8000}\times 100=37.5"

"\frac{37.5}{7}\times12=64.3"